OpenStudy (anonymous):

What is the exact value for the expression the square root of 56. − the square root of 14. + the square root of 126.? Simplify if possible. the square root of 14. 4 the square root of 14. 2 the square root of 42. 8 the square root of 42.

OpenStudy (anonymous):

sqr root of 56 - sqr root of 14 + sqr root of 126 = sqr root of (4x14) - sqr root of 14 + sqr root of (9x14) = 2(sqr root of 14) - (sqr root of 14) + 3(sqr root of 14) = 4(sqr root of 14)